OpenStudy (anonymous):

The product of (a + b)(a − b) is a2 − b2. Sometimes Always Never

OpenStudy (ash2326):

@SOS101 Can you multiply these? \[(a+b)\times (a-b)\]

OpenStudy (anonymous):

no because they are only variables , am i right?

OpenStudy (ash2326):

we can multiply variables \[x\times x=x^2\] \[x\times y = xy\]

OpenStudy (anonymous):

Oh lol sorry im not to good with these things

OpenStudy (anonymous):

So a^2 and b^2

OpenStudy (anonymous):

So would this be always @ash2326

OpenStudy (ash2326):

Could you multiply and check, what would you get?

OpenStudy (anonymous):

a^2 and b^2 but im not sure if i subtract them

OpenStudy (anonymous):

@SA5UK3

OpenStudy (anonymous):

Im not entirley getting the question

OpenStudy (ash2326):

Let's multiply and check \[(a+b)\times (a-b)\] \[=>\] \[a\times (a-b)+b\times (a-b)\] Do you agree with this?

OpenStudy (anonymous):

Its a difference of squares.

OpenStudy (anonymous):

So its always true

OpenStudy (anonymous):

Yes I agree and @SA5UK3 Ohhhh